# BACHI: Boundary-Aware Symbolic Chord Recognition Through Masked Iterative Decoding on Pop and Classical Music

### v1

@@ 2.1 Input and Output
For input, BACHI encodes symbolic musical scores as piano rolls $\mathbf{P}\in\{0,1\}^{T\times D}$, where $D=88$ is the dimensionality of pitch classes and the temporal resolution for $T$ is 12 frames per beat. We employ a patch embedding module to convert piano rolls into continuous latent tokens. Specifically, the patch embedding module contains a 1D-CNN layer with a kernel size 6 to map the pitch class channel $D$ to the hidden dimension of the transformer d_model=512d\_model=512, and reduce the temporal dimension from $T$ to $T/6$. It is followed with a GLU activation layer to create a more normalized and compact representation.
@@ 2.2 Two-stage Chord Recognition
At inference, we follow the below steps to obtain the final chord prediction:
@@ 3.2 Evaluation, Ablations and Baselines
We also conduct ablation studies on four variants: (1) BACHI without boundary detection and iterative decoding, directly projecting encoder outputs to chord labels; (2) BACHI with hard boundary constraints; (3) BACHI conditioned on additional key detection, implemented by training an auxiliary MLP for key prediction and embedding its output via FiLM together with boundary information; and (4) the full BACHI model. Due to page limitations, we report results on classical datasets in the main text, with POP909-CL results provided in the supplementary website. [^1] https://andyweasley2004.github.io/BACHI/.
@@ 4.1 Comprehensive Performance
Table 1 presents results across both classical and pop music evaluation sets. For classical music, BACHI achieves the highest scores across all metrics, with a notable improvement in full chord accuracy (69%) over prior baselines. Nevertheless, chord recognition in classical music remains highly challenging, as all models perform the full chord accuracy below 70%. This reflects the greater harmonic complexity and stylistic biases across composers and periods, highlighting the need for future work to better address these challenges.
@@ 4.2 Ablation Study
Table 2 presents ablation results of BACHI variants on the classical corpus, where BD denotes boundary detection and ID denotes iterative decoding. Without BD and ID, BACHI reduces to a basic transformer encoder model, achieving reasonable individual component performance (78.4% root, 79.5% quality, 77.6% bass) but lower full chord accuracy (66.8%).
Including BD alone improves results, but the absence of iterative decoding still limits performance, indicating that iterative decoding across chord elements contributes significantly to overall accuracy. Interestingly, adding key detection as an additional conditioning signal slightly decreases full-chord accuracy compared to the full BACHI, likely due to errors in key prediction propagating to chord recognition.
@@ 6 Conclusion
Our boundary-aware framework and POP909-CL annotations provide a foundation for advancing generative systems, enabling more harmonically coherent and structurally aware generation under chord conditions. Beyond applications in generation, learned confidence patterns also offer new insights into music theory by highlighting repertoire-specific harmonic tendencies. We will release our code, trained models, and POP909-CL annotations soon.

### v2

@@ 2.1 Input and Output
For input, BACHI encodes symbolic musical scores as piano rolls $\mathbf{P}\in\{0,1\}^{T\times D}$, where $D=88$ is the dimensionality of pitch classes and the temporal resolution for $T$ is 12 frames per beat, keeping triplet and dotted 16th notes precise. We employ a patch embedding module to convert piano rolls into continuous latent tokens. Specifically, the patch embedding module contains a 1D-CNN layer with a kernel size 6 to map the pitch class channel $D$ to the hidden dimension of the transformer d_model=512d\_model=512, and reduce the temporal dimension from $T$ to $T/6$. It is followed with a GLU activation layer to create a more normalized and compact representation.
@@ 2.2 Two-stage Chord Recognition
At inference, we follow the steps below to obtain the final chord prediction:
@@ 3.2 Evaluation, Ablations and Baselines
We also conduct ablation studies on four variants: (1) BACHI without boundary detection and iterative decoding, directly projecting encoder outputs to chord labels; (2) BACHI without iterative decoding; (3) BACHI conditioned on additional key detection, implemented by training an auxiliary MLP for key prediction and embedding its output via FiLM together with boundary information; and (4) the full BACHI model. Due to page limitations, we report all accuracy comparison and ablation results on classical datasets in the main text, with POP909-CL results and full MIREX-style comparisons provided in the demo website. [^1] https://andyweasley2004.github.io/BACHI/.
@@ 4.1 Comprehensive Performance
Table 1 presents results across both classical and pop music evaluation sets. For classical music, BACHI achieves the highest scores across all metrics, with a notable improvement in full chord accuracy (68%) over prior baselines. Nevertheless, chord recognition in classical music remains highly challenging, as all models perform the full chord accuracy below 70%. This reflects the greater harmonic complexity and stylistic biases across composers and periods, highlighting the need for future work to better address these challenges.
@@ 4.2 Ablation Study
Table 2 presents ablation results of BACHI variants on the classical corpus, where BD denotes boundary detection and ID denotes iterative decoding. Without BD and ID, BACHI reduces to a basic transformer encoder model, achieving reasonable individual component performance (78.0% root, 78.9% quality, 77.3% bass) but lower full chord accuracy (66.8%).
Adding BD primarily helps by reducing jitter (in our results, this gain is most evident on pop). However, the absence of iterative decoding still limits performance, indicating that iterative decoding across chord elements contributes significantly to overall accuracy. Interestingly, adding key detection as an additional condition slightly decreases full-chord accuracy compared to the full BACHI, likely due to errors in key prediction propagating to chord recognition.
@@ 6 Conclusion
Our boundary-aware framework and POP909-CL annotations provide a foundation for advancing generative systems, enabling more harmonically coherent and structurally aware generation under chord conditions. Beyond applications in generation, learned confidence patterns also offer new insights into music theory by highlighting repertoire-specific harmonic tendencies. We have released our code, trained models, and POP909-CL annotations, which can be accessed on our demo page  [^2] https://andyweasley2004.github.io/BACHI/.